OpenStudy (anonymous):

The equation s=p(1+r)tcan be used to model a person’s future income. S = future salary, p = current salary, r = rate of increase, and t = time in years. Susan’s salary increases by 3% each year. If she starts at $43,000 per year, what will her salary be in 5 years?

OpenStudy (anonymous):

s = p(1+r)^t note the exponential form s = 43000(1.03)^5